OpenStudy (anonymous):

what expression is equivalent to (m-^9) (-4 m^ - 5) (- m ^ 6)

OpenStudy (jdoe0001):

\(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\\ \quad \\ \quad \\ (m^{-{\color{red}{ 9}}})(-4m^{-{\color{red}{ 5}}})(-m^6)\implies \left(\cfrac{1}{m^{{\color{red}{ 9}}}}\right)\left(-4\cdot\cfrac{1}{m^{{\color{red}{ 5}}}}\ \right)(-m^6) \\ \quad \\ \left(\cfrac{1}{m^{{\color{red}{ 9}}}}\right)\left(\cfrac{-4}{m^{{\color{red}{ 5}}}}\ \right)(-m^6)\) multiply and simplify, see what you get